Construct the Confidence Interval for p if q = .364 and n = 2,200. (Use 1.96)
Answer #1

Solution :

Given that,

n = 2200

q=0.364

Point estimate = sample proportion = \hat p = 1-q=1-0.364=0.636

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96   ( Using z table )

Margin of error = E = Z\alpha / 2    * \sqrt (((\hat p * (1 - \hat p )) / n)

= 1.96 (\sqrt((0.636*0.364) /2200 )

= 0.0201

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.636-0.0201 < p <0.636+ 0.0201

0.6159< p < 0.6561

The 95% confidence interval for the population proportion p is : 0.6159,0.6561